QUESTION 1

A certain large manufacturing facility produces 20,000 parts each week. The manager of the facility estimates that about 1% of the parts they make are defective. What is the variance for the number of defective parts made each week?
1
Expert's answer
2020-02-24T11:28:22-0500

This is a case of binomial distribution with n=20000, p=0.01,q=1-p, 1-0.01=0.99

Variance=npq

=20000*0.01*0.99

=198